# A model for Axial Non-Standard Interactions of neutrinos with quarks

### v1

@@ II The model
Let us now discuss whether $\chi_{L}$ can play the role of Dark Matter (DM). To cancel the $U(1)-U(1)-U(1)$ anomaly, we had to add a SM singlet chiral field, $\chi_{L}$. In principle, our gauge group allows a coupling of form $\bar{\chi_{L}^{c}}H^{T}cL_{\tau}$ which leads to the decay of $\chi_{L}$ to $\nu_{\tau}$ plus muon or $s$-quarks pairs. Imposing a $Z_{2}$ symmetry under which only $\chi_{L}$ is odd this term will be forbidden and $\chi_{L}$ can be stable. Integrating out the intermediate $Z^{\prime}$, the effective coupling between $\chi$ and the first generation quarks can be written as
$\frac{9g_{Z}^{2}}{m_{Z^{\prime}}^{2}}\left(\bar{\chi}\gamma^{\mu}(\frac{1-\gamma_{5}}{2})\chi\right)\left(\bar{q}\gamma_{\mu}\gamma_{5}q\right).$ (16)
@@
$\sigma_{SD}^{N}\simeq 20.97\left(\frac{m_{N}m_{\chi}}{m_{N}+m_{\chi}}\right)^{2}\frac{g_{Z}^{4}}{m_{Z^{\prime}}^{4}}=5.24\left(\frac{m_{N}m_{\chi}}{m_{N}+m_{\chi}}\right)^{2}\frac{1}{\Big{(}(2M_{T}/\lambda_{T})^{2}+(18m_{\chi}/\lambda_{\chi})^{2}\Big{)}^{2}}\ .$ (17)
Notice that our model is isospin invariant, predicting equal cross sections for scattering off the neutron and the proton. Fig 2 shows the prediction of our model for the scattering cross section for different values of $\lambda_{T}=\lambda_{B}$ along with the bounds from the spin dependent direct detection. As seen from these figures, the direct dark matter detection bounds from the PICO experiment Amole et al. (2019) forces $\chi$ to be lighter than a few GeV:
@@
Taking $m_{Z^{\prime}}>30$ GeV, we can therefore use the four Fermi effective formalism to compute the cross section of the annihilation of dark matter pair to the standard model fermions Beltran et al. (2009). The cross section of the annihilation to the SM fermions for $\langle\phi\rangle>400$ GeV (i.e., $M_{T},M_{B}>1600$ GeV) will be much smaller than 1 pb, satisfying the bounds from the indirect dark matter searches Beltran et al. (2009). However, if the reheating temperature is higher than the mass of $\chi$ and $Z^{\prime}$, they will come into thermal equilibrium with the plasma in the early universe. An over-abundance of DM can be avoided if the reheating temperature is low Boddy et al. (2024). To have freeze-out mechanism with the annihilation cross section $\sigma_{ann}\sim 1$ pb, we need to turn on new interaction(s). The $\phi^{\prime}$ particle can be a $s$-channel mediator for this annihilation. For example, if $\phi^{\prime}$ mixes with the neutral component of a $SU(2)$ triplet with the same $U(1)$ charge, we can obtain annihilation into neutrinos, $\chi\chi\to\nu_{\tau}\nu_{\tau}$, which is less constrained by indirect dark matter searches than annihilation to quarks or tau. Another option is a coupling of $\phi^{\prime}$ to two light singlet scalars, $a$. This can take place via a quartic coupling of form $\lambda_{a\phi^{\prime}}|\phi^{\prime}|^{2}a^{2}$. The annihilation cross section will be given by
$\sigma v=\frac{\lambda_{a\phi^{\prime}}^{2}m_{\chi}^{2}}{128\pi(m_{\phi^{\prime}}^{2}-4m_{\chi}^{2})^{2}}\sqrt{1-\frac{m_{a}^{2}}{m_{\chi}^{2}}}\Big{(}1+\mathcal{O}(v^{2})\Big{)},$ (19)
which can be equal to 1 pb for $m_{\phi^{\prime}}\sqrt{1-4m_{\chi}^{2}/m_{\phi^{\prime}}^{2}}=31~{}{\rm GeV}\sqrt{\lambda_{a\phi^{\prime}}m_{\chi}/{\rm GeV}}$. If $m_{\chi}\sim$GeV, the freeze-out will take place around the time of QCD phase transition so the density of $a$ at the neutrino decoupling era will be suppressed and it will contribute less than 1 to the relativistic degrees of freedom ($\Delta N_{eff}$) at the neutrino decoupling era and later. At temperatures below $m_{\chi}/20$, the $a$ particles will decouple from the plasma and their density will be entropy suppressed. Taking the $a$ mass smaller than $\sim 100$ eV, their contribution to the matter budget of the universe today will be negligible, making $\chi_{L}$ the prime dark matter candidate.
@@
The $\chi$ particle can scatter off each other via $Z^{\prime}$ and/or $\phi^{\prime}$ exchange but the cross section will be much below values that can have observable effects Beltran et al. (2009): $\sigma(\chi\chi\to\chi\chi)/m_{\chi}\ll 0.1~{}{\rm cm}^{2}/{\rm gr}.$ As a result, our dark matter candidate can be regarded as collision-less dark matter for structure formation and galaxy cluster collisions.
@@
Figure 2: Spin-dependent DM-nucleon scattering cross section vs. DM mass for (a) protons and (b) neutrons. The solid curves are the present bounds from spin dependent direct dark matter search experiments and the dashed curves are the predictions of our model. The magenta solid line represents PICO-60 Amole et al. (2019), the red solid line shows LZ 2022 Aalbers et al. (2023), the green solid line depicts XENONnT 2023 Aprile et al. (2023), the brown solid line demonstrates PandaX-II Xia et al. (2019), the orange dashed line represents LUX-2017 Aalbers et al. (2023), and the cyan solid line shows LUX-2013 Akerib et al. (2016). The spin dependent cross section $\sigma^{N}_{SD}$ is plotted for $\lambda_{\chi}=1$ and four different values: $\lambda_{B}=\lambda_{T}=4$ (dashed blue), $\lambda_{B}=\lambda_{T}=3$ (dashed purple), $\lambda_{B}=\lambda_{T}=2$ (dashed pink), and $\lambda_{B}=\lambda_{T}=1$ (dashed gray), fixing $M_{T}=M_{B}=1550$ GeV.
@@ III Testable signals of the model
In the previous section, we found that in our model the axial NSI coupling, $\epsilon_{\tau\tau}^{Au}=\epsilon_{\tau\tau}^{Ad}$, can be as large as $U(1)$ which as shown in Abbaslu et al. (2024) will have observable impact on the number of neutral current events at the far detector of DUNE. This coupling can also affect the diffusion of the tau neutrinos out of the supernova core Cerdeño et al. (2023); Fiorillo et al. (2024); Cerdeño et al. (2021). In fact, the main goal of the present model is to obtain axial NSI with large couplings. To build a viable anomaly free gauge interaction that leads to such NSI, we introduce new particles, $Z^{\prime}$, heavy quarks ($T$ and $B$) and the dark matter candidate ($\chi_{L}$). An exciting aspect of our model is that each of these particles may show up in the upcoming experiments, making our toy model “super-testable.” Let us discuss their potential signatures one by one.
@@
Discovery of $Z^{\prime}$: A low mass gauge boson such as $Z^{\prime}$ coupled to $u$ and $d$ can be discovered as a resonance in the dijet distribution in the proton proton colliders. In fact, the LHC data with an integrated luminosity of 35.9 fb-1 has already set a bound of $\sim 0.2$ on the coupling of $Z^{\prime}$ with a mass of $\sim 30$ GeV to the quarks Sirunyan et al. (2019). The high luminosity phase of the LHC with 3 ab-1 integrated luminosity (100 times more data) is expected to probe couplings one order of magnitude lower, delving into the most interesting range of the parameter space of this model. However, this discovery cannot determine whether the $Z^{\prime}$ coupling to the quarks is vectorial or axial. Observing a deviation of the number of the neutral current events at the far detector of DUNE, along with the stringent bounds on $\epsilon^{Vq}$ from the other neutrino experiments, can testify the axial form of the interaction.
Signatures of $\chi_{L}$: As we saw in the previous section, this model naturally offers a candidate for a dark matter which can undergo spin dependent dark matter scattering off nuclei. Direct dark matter search experiments already constrain dark matter mass to be lighter than $O(1)$ GeV within this model with $\epsilon^{Au}=\epsilon^{Ad}\sim 1$. The upcoming direct dark matter searches can find $\chi_{L}$. The indirect dark matter search bounds from dwarf galaxies are satisfied in our model. If the production mechanism for $\chi_{L}$ in the early universe is freeze-out with annihilation to dark sector, we would expect fractional contribution to extra relativistic degrees of freedom in the early universe which can be eventually tested by improvements on the BBN and CMB measurements.
@@
Discovery of $T$ and $B$: As emphasized in the previous section $t_{R}$, $c_{R}$, $b_{R}$ and $s_{R}$ with $U(1)$ charge of 1 are taken to be mass eigenstates. However, mass eigenstates $\tilde{u}_{R}$ and $\tilde{T}_{R}$ ($\tilde{d}_{R}$ and $\tilde{B}_{R}$) are a mixture of $u_{R}$ and $T_{R}$ ($d_{R}$ and $B_{R}$) with opposite $U(1)$ charges. The same consideration applies for left-handed components of these quarks, $\tilde{u}_{L}$ and $\tilde{T}_{L}$ ($\tilde{d}_{L}$ and $\tilde{B}_{L}$). In fact, in the case of left-handed components, we should also consider that while $u_{L}$ and $d_{L}$ form a $SU(2)$ doublet, $T_{L}$ and $B_{L}$ do not. Putting these considerations together, we find out that the flavor changing neutral currents with vertices including third and/or fourth generations are forbidden, implying that Kaons, $D$-mesons and $B$-mesons will behave as the SM predicts but we can have vertices such as
@@
given by the relevant gauge coupling times the corresponding mixing angle. Similarly, we have vertices
@@
where $\sin\alpha\simeq\lambda_{Tu}\langle H\rangle/M_{T}$ and $\sin\beta\simeq\lambda_{Bd}\langle H\rangle/M_{B}$. As mentioned in sect. II, the dominant decay modes of $\tilde{B}$ and $\tilde{T}$ are $\tilde{B}\to W\tilde{u}$, $\tilde{B}\to Z\tilde{d}$, $\tilde{T}\to W\tilde{d}$ and $\tilde{T}\to Z\tilde{u}$. Remember that $\epsilon^{Au}_{\tau\tau}=\epsilon^{Ad}_{\tau\tau}$ is suppressed by the inverse of $\langle\phi\rangle^{2}=M_{T}^{2}/\lambda_{T}^{2}=M_{B}^{2}/\lambda_{B}^{2}$. Thus, to have sizable NSI, the masses of $T$ and $B$ should be close to the present lower bound. $B$ and $T$ can therefore be pair produced via QCD interaction at the LHC and detected as jet+$W$ or jet+$Z$. As a result, there is a good chance of discovery of these particles in near future.

### v2

@@ II The model
Inserting $m_{u}\sim m_{d}\sim 5$ MeV in Eq. 10, we find that $\lvert\lambda_{Tu}M_{Tu}^{\star}\lvert\sim 0.045~{\rm GeV}(M_{T}/1550~{\rm GeV})$ and $\lvert\lambda_{Bd}M_{Bd}^{\star}\lvert\sim 0.045~{\rm GeV}(M_{B}/1550~{\rm GeV})$. $Q_{1}$ and $T_{R}$, being heavier than the Higgs, cannot appear in the decay modes of $H$. Moreover, we can take $\lambda_{Tu},\lambda_{Bd}\ll\lambda_{t}$ so that the radiative correction due to these couplings to Higgs observables such as the Higgs mass or the Higgs decay modes will be negligible compared to those from the SM top Yukawa coupling.
The Lagrangian can also include the following quartic terms
$\lambda_{H}|H|^{4}+\lambda_{\phi}|\phi|^{4}+\lambda_{\phi^{\prime}}|\phi^{\prime}|^{4}+\lambda_{\phi\phi^{\prime}}|\phi|^{2}|\phi^{\prime}|^{2}+\lambda_{H\phi}|H|^{2}|\phi|^{2}+\lambda_{H\phi^{\prime}}|H|^{2}|\phi^{\prime}|^{2}\ .$ (17)
The “unbounded from below condition” implies $\lambda_{H},\lambda_{\phi},\lambda_{\phi^{\prime}}>0$, $\lambda_{\phi\phi^{\prime}}>-2\sqrt{\lambda_{\phi}\lambda_{\phi^{\prime}}}$, $\lambda_{H\phi}>-2\sqrt{\lambda_{\phi}\lambda_{H}}$ and $\lambda_{H\phi^{\prime}}>-2\sqrt{\lambda_{\phi^{\prime}}\lambda_{H}}$. The couplings $\lambda_{H}$, $\lambda_{\phi}$ and $\lambda_{\phi^{\prime}}$ should be nonzero to induce vacuum expectation value for the corresponding scalar fields. However, $\lambda_{\phi\phi^{\prime}}$, $\lambda_{H\phi}$ and $\lambda_{H\phi^{\prime}}$ can be arbitrarily small or even zero without affecting the results of our paper. Indeed from various considerations, there are bounds on these couplings: The stability of masses of the scalars requires $\lambda_{\phi\phi^{\prime}}\ll m_{\phi^{\prime}}^{2}/\langle\phi\rangle^{2}$, $\lambda_{H\phi^{\prime}}\ll m_{\phi^{\prime}}^{2}/\langle H\rangle^{2}$ and $\lambda_{H\phi}\ll m_{H}^{2}/\langle\phi\rangle^{2}$. Moreover, to make $Br(H\to\phi^{\prime}\phi^{\prime})$ smaller than $Br(H\to b\bar{b})$, it is required $\lambda_{H\phi^{\prime}}\ll m_{b}/\langle H\rangle=0.02$. Throughout our analysis, we set $\lambda_{H\phi^{\prime}}=\lambda_{\phi\phi^{\prime}}=\lambda_{H\phi}=0$. Then, we simply have
$M_{H}=2\sqrt{\lambda_{H}}\langle H\rangle,\ m_{\phi}=2\sqrt{\lambda_{\phi}}\langle\phi\rangle\ \ {\rm and}\ \ m_{\phi^{\prime}}=2\sqrt{\lambda_{\phi^{\prime}}}\langle\phi^{\prime}\rangle\ .$ (18)
Let us now discuss whether $\chi_{L}$ can play the role of Dark Matter (DM). To cancel the $U(1)U(1)$-U(1)-U(1) anomaly, we had to add a SM singlet chiral field, $\chi_{L}$. In principle, our gauge group allows a coupling of form $\bar{\chi_{L}^{c}}\epsilon_{ij}H_{i}(L_{\tau})_{j}$ which leads to the decay of $\chi_{L}$ to $\nu_{\tau}$ plus muon or $s$-quarks pairs. Imposing a $Z_{2}$ symmetry under which only $\chi_{L}$ is odd this term will be forbidden and $\chi_{L}$ can be stable. Integrating out the intermediate $Z^{\prime}$, the effective coupling between $\chi$ and the first generation quarks can be written as
$\frac{9g_{Z^{\prime}}^{2}}{m_{Z^{\prime}}^{2}}\left(\bar{\chi}\gamma^{\mu}(\frac{1-\gamma_{5}}{2})\chi\right)\left(\bar{q}\gamma_{\mu}\gamma_{5}q\right).$ (19)
@@
$\sigma_{SD}^{N}\simeq 20.97\left(\frac{m_{N}m_{\chi}}{m_{N}+m_{\chi}}\right)^{2}\frac{g_{Z^{\prime}}^{4}}{m_{Z^{\prime}}^{4}}=5.24\left(\frac{m_{N}m_{\chi}}{m_{N}+m_{\chi}}\right)^{2}\frac{1}{\Big{(}(2M_{T}/\lambda_{T})^{2}+(18m_{\chi}/\lambda_{\chi})^{2}\Big{)}^{2}}\ .$ (20)
Notice that our model is isospin invariant, predicting equal cross sections for scattering off the neutron and the proton. Figure 2 shows the prediction of our model for the scattering cross section for different values of $\lambda_{T}=\lambda_{B}$ along with the bounds from the spin dependent direct detection. As seen from these figures, the direct dark matter detection bounds from the PICO experiment Amole et al. (2019) forces $\chi$ to be lighter than a few GeV:
@@
Notice that $\chi$ can still be much heavier than few keV which is the lower bound from the structure formation consideration Cirelli et al. (2024). Taking $m_{Z^{\prime}}>30$ GeV, we can therefore use the four Fermi effective formalism to compute the cross section of the annihilation of dark matter pair to the standard model fermions Beltran et al. (2009). The cross section of the annihilation to the SM fermions for $\langle\phi\rangle>400$ GeV (i.e., $M_{T},M_{B}>1600$ GeV) will be much smaller than 1 pb, satisfying the bounds from the indirect dark matter searches Beltran et al. (2009).
The aim of this paper is to build, for the first time, an underlying model for $\epsilon^{Au}_{\tau\tau}=\epsilon^{Ad}_{\tau\tau}\sim 1$. We realized that the model includes a colorless electroweak singlet which can be a suitable dark matter candidate. In the literature, a myriad of production mechanisms for dark matter has been suggested. Below, we briefly mention a few possibilities to be incorporated into the model. We emphasize that these mechanisms do not affect the phenomenology of neutrino interaction with matter which is the focus of the present paper. Being only peripheral to our discussion, we shall not go into details and will keep the focus on $\epsilon^{Au}=\epsilon^{Ad}$.
  • i) Low reheating temperature cosmologies: As shown in Ref. Hannestad (2004), the reheating temperature can be as low as 4 MeV. If the reheating temperature is smaller than the mass of DM, the initial DM number density would be Boltzmann suppressed. Recently a new class of DM production mechanisms has been suggested that is based on low reheating temperatures Gan and Tsai (2023); Boddy et al. (2024). This mechanism can be readily incorporated into our model without affecting our main results.
  • ii) Freeze-out scenario: As is well-known in order to obtain the observed DM abundance via the freeze-out scenario, the annihilation cross section should be of the order of 1 pb. A detailed study of the freeze-out scenario for DM of GeV mass scale is carried out in Steigman et al. (2012). To obtain $\sigma_{ann}\sim 1$ pb, we need to turn on new interaction(s). The $\phi^{\prime}$ particle can be a $s$-channel mediator for this annihilation. For example, if $\phi^{\prime}$ mixes with the neutral component of a $SU(2)$ triplet with the same $U(1)$ charge, we can obtain annihilation into neutrinos, $\chi\chi\to\nu_{\tau}\nu_{\tau}$, which is less constrained by indirect dark matter searches than annihilation to quarks or tau. Another option is the annihilation of the DM pair to a pair of light dark sector particles via $s$-channel $\phi^{\prime}$ exchange. Further elaboration is beyond the scope of the present paper.
@@
The $\chi$ particles can scatter off each other via $Z^{\prime}$ and/or $\phi^{\prime}$ exchange but the cross section will be much below the values that can have observable effects Peter et al. (2013): $\sigma(\chi\chi\to\chi\chi)/m_{\chi}\ll 0.1~{\rm cm}^{2}/{\rm gr}.$ As a result, our dark matter candidate can be regarded as collision-less dark matter for structure formation and galaxy cluster collisions. For a review on the impact of self-interaction on structure formation, see Tulin and Yu (2018).
@@
Figure 2: Spin-dependent DM scattering cross section off protons (upper panel) and off neutrons (lower panel) versus DM mass. The solid curves are the present bounds from spin dependent direct dark matter search experiments and the dashed curves are the predictions of our model. The blue solid line shows PandaX-4T 2022 Huang et al. (2022), the green solid line depicts XENONnT 2023 Aprile et al. (2023), the black solid line represents PandaX-4T 2024 Abdukerim et al. (2025), the red solid line shows LZ 2022 Aalbers et al. (2023) and finally, the magenta solid line represents PICO-60 Amole et al. (2019). The spin dependent cross section $\sigma^{N}_{SD}$ is plotted for $\lambda_{\chi}=1$, $M_{T}=M_{B}=1550$ GeV and four different values of the quark couplings: $\lambda_{B}=\lambda_{T}=4$ (dashed blue), $\lambda_{B}=\lambda_{T}=3$ (dashed purple), $\lambda_{B}=\lambda_{T}=2$ (dashed pink) and $\lambda_{B}=\lambda_{T}=1$ (dashed gray).
Let us now discuss the impact of our model on the early universe, with special attention to the possibility of extra relativistic degrees of freedom that can affect Big Bang Nucleosynthesis (BBN) and/or CMB. The new gauge boson ($Z^{\prime}$) and the new scalar ($\phi$) as well as the new quarks ($T$ and $B$) are all heavier than GeV and decay long before the BBN era and the neutrino decoupling so they cannot affect these observables. The impact of $\chi_{L}$ and $\phi^{\prime}$ on the early universe cosmology depends on their production mechanism. Within the low reheating mechanism from the beginning, their densities will be too small to be relevant. Within the freeze-out scenario, $\chi_{L}$ and $\phi^{\prime}$ can be produced and thermalized in the early universe via the new gauge interaction. The freeze-out of $\chi_{L}$ via $\chi_{L}\chi_{L}\to\nu_{\tau}\nu_{\tau}$ would take place at $T\sim m_{\chi}/20$. As long as $m_{\chi}/20\ggfew$ MeV (i.e., neutrino decoupling era), the annihilation products thermalize with the plasma and cannot affect BBN and CMB, either. $\phi^{\prime}$ can quickly decay into the $\chi_{L}$ pair even before the $\chi_{L}$ freeze-out so it cannot affect BBN and CMB.
Due to the $Z^{\prime}$ coupling to $\nu_{\tau}$, the cross section of self-scattering of $\stackrel{{\scriptstyle(-)}}{{\nu}}_{\tau}\stackrel{{\scriptstyle(-)}}{{\nu}}_{\tau}\to\stackrel{{\scriptstyle(-)}}{{\nu}}_{\tau}\stackrel{{\scriptstyle(-)}}{{\nu}}_{\tau}$ can increase. However, this enhancement cannot significantly affect the structure formation because at the matter radiation equality era ($T\sim 2$ eV) when the structure formation becomes efficient, the rate of this scattering becomes much lower than the Hubble rate: $n_{\nu}\sigma(\nu_{\tau}\nu_{\tau}\to\nu_{\tau}\nu_{\tau})|_{T\simeq 2~{\rm eV}}\ll T^{2}/M_{Pl}^{*}|_{T\simeq 2~{\rm eV}}$ where $\sigma(\nu_{\tau}\nu_{\tau}\to\nu_{\tau}\nu_{\tau})\sim(9g_{Z^{\prime}})^{4}T^{2}/(4\pi m_{Z^{\prime}}^{4}).$
@@ III Testable signals of the model
In the previous section, we found that in our model the axial NSI coupling, $\epsilon_{\tau\tau}^{Au}=\epsilon_{\tau\tau}^{Ad}$, can be as large as $\mathcal{O}(1)$ which as shown in Abbaslu et al. (2024) will have observable impact on the number of neutral current events at the far detector of DUNE. This coupling can also affect the diffusion of the tau neutrinos out of the supernova core Cerdeño et al. (2023); Fiorillo et al. (2024); Cerdeño et al. (2021). In fact, the main goal of the present model is to obtain axial NSI with large couplings. To build a viable anomaly-free gauge interaction that leads to such NSI, we introduced a new gauge boson, $Z^{\prime}$, heavy quarks ($T$ and $B$) and the dark matter candidate ($\chi_{L}$). An exciting aspect of our model is that each of these particles may show up in the upcoming experiments, making our toy model “super-testable.” Let us discuss their potential signatures one by one.
@@
Discovery of $Z^{\prime}$: A low mass gauge boson such as $Z^{\prime}$ coupled to $u$ and $d$ can be discovered as a resonance in the dijet distribution in the proton proton colliders. In fact, the LHC data with an integrated luminosity of 35.9 fb-1 has already set a bound of $\sim 0.2$ on the coupling of $Z^{\prime}$ with a mass of $\sim 30$ GeV to the quarks Sirunyan et al. (2019). The high luminosity phase of the LHC with 3 ab-1 integrated luminosity (100 times more data) is expected to probe couplings one order of magnitude lower, delving into the most interesting range of the parameter space of this model. However, this discovery cannot determine whether the $Z^{\prime}$ coupling to the quarks is axial or of vector-type. Observing a deviation of the number of the neutral current events at the far detector of DUNE, along with the stringent bounds on $\epsilon^{Vq}$ from the other neutrino experiments, can testify the axial form of the interaction.
Signatures of $\chi_{L}$: As we saw in the previous section, this model naturally offers a candidate for a dark matter which can undergo spin dependent dark matter scattering off nuclei. Direct dark matter search experiments already constrain the dark matter mass to the values lighter than $O(1)$ GeV within this model with $\epsilon^{Au}=\epsilon^{Ad}\sim 1$. The upcoming direct dark matter searches may find $\chi_{L}$. For the detection of a GeV scale DM, the main obstacle is lowering the detection energy threshold, $E_{th}$. The minimum velocity that a DM particle should have in order to lead to an observable recoil energy is $v_{min}\simeq\sqrt{2E_{th}/m_{DM}}$ Cirelli et al. (2024). The number of DM scattering events will be given by $\sigma_{SD}^{N}\times\int_{v_{min}}^{v_{esc}}\exp(-v^{2}/v_{0}^{2})v^{2}dv$ where $v_{0}\simeq 250$ km/sec and $v_{esc}\simeq 544$ km/sec Cirelli et al. (2024). As seen from Fig. 2, for $\epsilon_{\tau\tau}^{Au}=\epsilon_{\tau\tau}^{Ad}=0.9$ (i.e., $\lambda_{T}=\lambda_{B}=4$), the PICO experiment limits $m_{\chi}<2$ GeV. This experiment has an energy threshold of 2.45 keV Amole et al. (2019). If $E_{th}$ is lowered to 1.2 keV, with the same exposure, the number of events for $m_{DM}=1$ GeV will be the same as the number of events for $m_{DM}=2$ GeV and $E_{th}=2.45$ keV. That is $m_{DM}$ down to 1 GeV can be probed, covering the most interesting part of the parameter space.
@@
Discovery of $T$ and $B$: As emphasized in the previous section $t_{R}$, $c_{R}$, $b_{R}$ and $s_{R}$ with $U(1)$ charge of 1 are taken to be mass eigenstates. However, mass eigenstates $\tilde{u}_{R}$ and $\tilde{T}_{R}$ ($\tilde{d}_{R}$ and $\tilde{B}_{R}$) are a mixture of $u_{R}$ and $T_{R}$ ($d_{R}$ and $B_{R}$) with opposite $U(1)$ charges. The same consideration applies to left-handed components of these quarks, $\tilde{u}_{L}$ and $\tilde{T}_{L}$ ($\tilde{d}_{L}$ and $\tilde{B}_{L}$). In fact, in the case of left-handed components, we should also consider that while $u_{L}$ and $d_{L}$ form a $SU(2)$ doublet, $T_{L}$ and $B_{L}$ do not. Putting these considerations together, we find out that the flavor changing neutral currents with vertices including third and/or fourth generations are forbidden, implying that Kaons, $D$-mesons and $B$-mesons will behave as the SM predicts. However, we can have vertices such as
@@
given by the relevant gauge coupling times the corresponding mixing angle. Similarly, we have the following vertices
@@
where $\sin\alpha\simeq\lambda_{Tu}\langle H\rangle/M_{T}$ and $\sin\beta\simeq\lambda_{Bd}\langle H\rangle/M_{B}$. As mentioned in sect. II, the dominant decay modes of $\tilde{B}$ and $\tilde{T}$ are $\tilde{B}\to W\tilde{u}$, $\tilde{B}\to Z\tilde{d}$, $\tilde{T}\to W\tilde{d}$ and $\tilde{T}\to Z\tilde{u}$. Remember that $\epsilon^{Au}_{\tau\tau}=\epsilon^{Ad}_{\tau\tau}$ is suppressed by the inverse of $\langle\phi\rangle^{2}=M_{T}^{2}/\lambda_{T}^{2}=M_{B}^{2}/\lambda_{B}^{2}$. Thus, to have sizable NSI, the masses of $T$ and $B$ should be close to the present lower bounds. $B$ and $T$ can therefore be pair produced via QCD interaction at the LHC and detected as jet+$W$ or jet+$Z$. With 140 fb-1, the ATLAS collaboration has ruled out $T$ and $B$ quarks of 1530 GeV mass Aad et al. (2024b). That is the ratio of signal to root of background is equal or less than 2 (i.e., $\mathcal{S}/\sqrt{\mathcal{B}}\leq 2$). If the masses of these particles are close to 1500 GeV, this means that with 140 $fb^{-1}$ integrated luminosity, the ratio $\mathcal{S}/\sqrt{\mathcal{B}}\simeq 2$. At the high luminosity phase of LHC with 3 ab-1, both the signal and the background will be larger by a factor of $3~{\rm ab^{-1}}/140~{\rm ab^{-1}}=3000/140$. Thus, the ratio $\mathcal{S}/\sqrt{\mathcal{B}}$ can become as large as $\sqrt{3000/140}\times 2\sim 9.2$ which means there will be a good chance of $T$ and $B$ discovery at more than $5\sigma$ C.L.